### v1

# Inverse Virtual Try-On: Generating Multi-Category Product-Style Images from Clothed Individuals

Abstract
While virtual try-on (VTON) systems aim to render a garment onto a target person image, this paper tackles the novel task of virtual try-off (VTOFF), which addresses the inverse problem: generating standardized product images of garments from real-world photos of clothed individuals. Unlike VTON, which must resolve diverse pose and style variations, VTOFF benefits from a consistent and well-defined output format – typically a flat, lay-down-style representation of the garment – making it a promising tool for data generation and dataset enhancement. However, existing VTOFF approaches face two major limitations: (i) difficulty in disentangling garment features from occlusions and complex poses, often leading to visual artifacts, and (ii) restricted applicability to single-category garments (e.g., upper-body clothes only), limiting generalization. To address these challenges, we present Text-Enhanced MUlti-category Virtual Try-Off (TEMU-VTOFF), a novel architecture featuring a dual DiT-based backbone with a modified multimodal attention mechanism for robust garment feature extraction. Our architecture is designed to receive garment information from multiple modalities like images, text, and masks to work in a multi-category setting. Finally, we propose an additional alignment module to further refine the generated visual details. Experiments on VITON-HD and Dress Code datasets show that TEMU-VTOFF sets a new state-of-the-art on the VTOFF task, significantly improving both visual quality and fidelity to the target garments. Our project page is available at https://temu-vtoff-page.github.io/.
Figure 1: Visual results produced by our proposed text-enhanced multi-category virtual try-off architecture, i.e., TEMU-VTOFF. Given a clothed input person image, the proposed model reconstructs the clean, in-shop version of the worn garment. Our model handles various garment types and preserves both structural fidelity and fine-grained textures, even under occlusions and complex poses, thanks to its multimodal attention and garment-alignment design.
1 Introduction
Unlike virtual try-on (VTON), whose goal is to dress a given clothing image on a target person image, in this paper, we focus exactly on the opposite, virtual try-off (VTOFF), whose purpose is to generate standardized product images from real-world clothed individual photos. Compared to VTON, which often struggles with the ambiguity and diversity of valid outputs, such as stylistic variations in how a garment is worn, VTOFF benefits from a clearer output objective: reconstructing a consistent, lay-down-style image of the garment. This reversed formulation facilitates a more objective evaluation of garment reconstruction quality.
The fashion industry, a trillion-dollar global market, is increasingly integrating AI and computer vision to optimize product workflows and enhance user experience. VTOFF, in this context, offers substantial value: it enables the automatic generation of tiled product views, which are essential for tasks such as image retrieval, outfit recommendation, and virtual shopping. However, acquiring such lay-down images is expensive and time-consuming for retailers. VTOFF provides a scalable alternative by leveraging images of garments worn by models or customers, transforming them into standardized catalog views through image-to-image translation techniques.
Despite the recent success of GANs goodfellow2014GAN and Latent Diffusion Models (LDMs) rombach2022ldm in image translation tasks siarohin2019first ; ren2023pi ; isola2017image ; tumanyan2023plug , current VTOFF solutions face notable limitations. Existing models velioglu2024tryoffdiff ; xarchakos2024tryoffanyone struggle to accurately reconstruct catalog images from dressed human inputs. This limitation arises from a fundamental architectural mismatch – these approaches repurpose VTON pipelines by merely reversing the input-output roles, without addressing the unique challenges of the VTOFF task. Moreover, the high visual variability of real-world images – due to garment wear category (e.g., upper-body), pose changes, and occlusions – makes it difficult for these models to robustly extract garment features while preserving fine-grained patterns. On the opposite side, we design a dedicated architecture tailored for the VTOFF task, accounting for the unique challenges of complex positions and occlusions in person images as opposed to a flat catalog garment.
Recent advances in diffusion models demonstrate that DiT-based architectures peebles2023scalable , especially when combined with flow-matching objectives lipman2022flow , surpass traditional U-Net and DDPM-based approaches rombach2022ldm . Inspired by these findings, we propose TEMU-VTOFF, a Text-Enhanced MUlti-category Virtual Try-Off architecture based on a dual-DiT framework. Specifically, we exploit the representational strength of DiT in two distinct ways: (i) the first Transformer component focuses on extracting fine-grained garment features from complex, detail-rich person images; and (ii) the second DiT is specialized for generating the clean, in-shop version of the garment. To support this design, we further adapt the base DiT architecture to accommodate the task-specific input modalities. To further enhance alignment, we introduce an external garment aligner module and a novel supervision loss that leverages clean garment references as guidance, further improving quality of generated images.
Our contribution can be summarized as follows:
  • Multi-Category Try-Off. We present a unified framework capable of handling multiple garment types (upper-body, lower-body, and full-body clothes) without requiring category-specific pipelines.
  • Multimodal Hybrid Attention. We introduce a novel attention mechanism that integrates garment textual descriptions into the generative process by linking them with person-specific features. This helps the model synthesize occluded or ambiguous garment regions more accurately.
  • Garment Aligner Module. We design a lightweight aligner that conditions generation on clean garment images, replacing conventional denoising objectives. This leads to better alignment consistency on the overall dataset and preserves more precise visual retention.
  • Extensive experiments on the Dress Code and VITON-HD datasets demonstrate that TEMU-VTOFF outperforms prior methods in both the quality of generated images and alignment with the target garment, highlighting its strong generalization capabilities.
2 Related Work
Virtual Try-On. As one of the most popular tasks within the fashion domain, VTON has been widely studied over the past decades by the computer vision and graphics communities due to its interesting research challenges and the practical potential bai2022single ; cui2021dressing ; fele2022c ; ren2023cloth ; ge2021disentangled . Existing methods are broadly categorized into warping-based chen2023size ; xie2023gpvton ; yan2023linking and warping-free approaches zhu2023tryondiffusion ; morelli2023ladi ; baldrati2023multimodal ; zeng2024cat ; chong2025catvton , with a growing shift from GAN-based goodfellow2020generative to diffusion-based ho2020denoising ; song2020denoising frameworks for better image fidelity and stability. Warping-based methods usually follow a two-stage pipeline: first warping the garment using TPS bookstein1989principal , flow, or landmarks to align it with the body, then fusing it with the person image. VITON han2018viton , CP-VTON wang2018toward , and their variants improve garment alignment and synthesis quality, but often produce artifacts due to imperfect warping. To mitigate this, warping-free methods leverage diffusion models to bypass explicit deformation zhu2023tryondiffusion ; morelli2023ladi ; xu2024ootdiffusion ; choi2024idmimproving employing modified cross-attention or self-attention to directly condition generation on garment features, often using CLIP-based encoders radford2021learning . However, these pre-trained encoders tend to lose fine-grained texture details, prompting methods like StableVITON kim2023stableviton to introduce dedicated garment encoders and attention mechanisms, albeit at a higher computational cost. Lately, DiT-based works jiang2024fitdit show the benefits of Transformer-based diffusion models for high-fidelity garment to person transfer. While most existing works focus on generating dressed images from separate garment and person inputs, the inverse problem (i.e., reconstructing clean, standalone garment representations from worn images) remains underexplored.
Virtual Try-Off. While VTON has been extensively studied for synthesizing images of a person wearing a target garment, the recently proposed VTOFF task shifts the focus toward garment-centric reconstruction, aiming to extract a clean, standardized image of a garment worn by a person. TryOffDiff velioglu2024tryoffdiff introduces this task by leveraging a diffusion-based model with SigLIP zhai2023sigmoid conditioning to recover high-fidelity garment images. Building on this direction, TryOffAnyone xarchakos2024tryoffanyone addresses the generation of tiled garment images from dressed photos for applications like outfit composition and retrieval. By integrating garment-specific masks and simplifying the Stable Diffusion pipeline through selective Transformer tuning, it achieves both quality and efficiency. In both cases, these works have been designed for single-category scenarios, thus limiting their potential application to generate wider, more diverse data collections. On a different line, Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition is not a native VTOFF method, but it leverages a LoRA-based module hu2022lora to fine-tune FLUX flux2024 for this task. Though these works collectively reflect a growing shift from person-centric synthesis to garment-centric understanding, there are still limitations like frequent garment structural artifacts (e.g., in shape, neckline, waist) and on colors and textures of generated outputs. We hypothesize that this mismatch is due to a too generic architectural choice, untailored for the specific needs of the VTOFF setting. In this work, we focus on existing VTOFF open problems, such as multi-category adaptation, occlusions, and complex human poses, and propose a novel VTOFF-specific architecture enhanced with text and fine-grained mask conditioning and optimized with a garment aligner component that can improve the quality of generated garments.
3 Methodology
Figure 2: Overview of our method. The feature extractor $F_{E}$ processes spatial inputs (noise, masked image, binary mask), and global inputs (model image via AdaLN). The intermediate keys and values $\bm{K}^{l}_{\text{extractor}}$,$\bm{V}^{l}_{\text{extractor}}$ are injected into the corresponding hybrid blocks of the garment generator $F_{D}$. Then, the main DiT model generates the final garment leveraging the proposed MHA module. We align our model with a diffusion loss for the noise estimate and an alignment loss with clean, DINOv2 features of the target garment.
Preliminary. The latest diffusion models are a family of generative architectures that work by corrupting a ground-truth image $\bm{z}_{0}$ following a flow-matching schedule lipman2022flow defined as
$\bm{z}_{t}=(1-t)\bm{z}_{0}+t\epsilon_{t}\quad\epsilon\sim\mathcal{N}(0,1),\quad t\in[0,1].$ (1)
Then, a diffusion model estimates back the injected noise $\epsilon_{t}$ through a Diffusion Transformer (DiT) peebles2023scalable , obtaining a prediction $\hat{\bm{z}_{0}}$. In Stable Diffusion 3 (SD3) esser2024scaling , the 16-channel latent $\bm{z}_{t}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 16}$ is obtained projecting the original RGB image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ with a variational autoencoder $\mathcal{E}$ kingma2013vae , obtaining $\bm{z}=\mathcal{E}(\bm{x})$, with $H,W$ being height and width of the image, and $f=8$ the spatial compression ratio of the autoencoder. Finally, the model is trained according to an MSE loss function $\mathcal{L}_{\text{diff}}$:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\bm{z}_{0},\epsilon_{t},t}\left[\left\|\epsilon_{t}-\epsilon_{\theta}(\bm{z}_{t},t)\right\|^{2}\right].$ (2)
Overview. An overview of our method is shown in Fig. 2. The purpose of this framework is to generate an in-shop version of the garment worn by the person. The critical design choice is to process correctly the dressed person image in order to extract meaningful information to be injected in the denoising process. For this reason, we encompass a dual-DiT architecture, based on SD3, whose models are deputed to two different purposes. Firstly, we design the first DiT as a feature extractor $F_{E}$ that encodes the model image $\bm{x}_{\text{model}}$ and outputs its intermediate layer features at timestep $t=0$ and not from subsequent timesteps, as we are interested in extracting clean features from $F_{E}$. This block is trained with a diffusion loss to generate the person image. Once trained, this model outputs meaningful key and value features of the dressed person. Secondly, the main DiT generates the garment $\bm{x}_{g}$ leveraging the intermediate features from $F_{E}$ in a modified textual-enhanced attention module. The two DiT architectures can be manipulated along three axes: (i) the Transformer projector before the DiT to account for additional inputs; (ii) the Modulation Space (of every DiT block) to account for different input modalities (such as text and images); (iii) the Attention operator (in all layers) for fine-grained conditioning.
3.1 DiT Feature Extractor
We design the feature extractor $F_{E}$ as a DiT, working with two different input types: a global input with the person image $\bm{x}_{model}\in\mathbb{R}^{H\times W\times 3}$ leveraged by the modulation layers of $F_{E}$, and a local spatial input as the channel-wise concatenation $\bm{z}_{t}^{\prime}=[\bm{z}_{t},M,\bm{x}_{M}]\in\mathbb{R}^{h\times w\times 33}$ of the latent $\bm{z}_{t}$, the encoded latent of the masked person image $\bm{x}_{M}=\mathcal{E}(I_{M})\in\mathbb{R}^{h\times w\times 16}$ and the interpolated binary mask $M\in\mathbb{R}^{h\times w\times 1}$ encoded through the Transformer projector $\mathcal{P}:\mathbb{R}^{h\times w\times 33}\rightarrow\mathbb{R}^{S\times d}$, with $S$ as sequence length and $d$ as embedding dimension. We train it as a single DiT separated from the garment generator $F_{D}$. We extract the keys and values $\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}$ from every attention block $A^{l}$ of $F_{E}$, with $l=1,\ldots,N$, as shown in Fig. 2, because they bring valuable visual information about the garment features in the person image.
Mask Conditioning with modified Transformer Projector. We investigate whether fine-grained clothing attributes – such as garment shape, edge refinement, and neckline structure – can be inferred from text alone. These attributes are inherently structural and spatial, which poses a challenge for language-based conditioning due to its inherently diffuse nature. We posit that a segmentation mask can act as a “hard” discriminator, providing precise spatial constraints that complement the “soft” and distributed cues present in text.
To this end, we propose to use the mask $M$ and the associated masked person image $\bm{\bm{x}}_{M}$ as an additional conditioning of the extractor $F_{E}$ and we get a new latent $\bm{z}_{t}^{\prime}=[\bm{z}_{t},M,\bm{x}_{M}]\in\mathbb{R}^{h\times w\times 33}$. We modify the original SD3 DiT projector $\mathcal{P_{\text{orig}}}:\mathbb{R}^{h\times w\times 16}\rightarrow\mathbb{R}^{S\times d}$ by adding zero-initialized convolutional layers on the channel dimension, thus accounting for the increased channel-size of our input $\bm{z}_{t}^{\prime}$ and obtaining the final projector $\mathcal{P}$. We train the module $F_{E}$ alone, detached from the dual DiT $F_{D}$, according to the diffusion loss $L_{\text{diff}}$ defined as follows:
$\mathcal{L}_{\text{extractor}}=\mathbb{E}_{\bm{z}_{0},\epsilon_{t},t}\left[\left\|\epsilon_{t}-F_{E}(\bm{z}_{t}^{\prime},\bm{x}_{\text{model}},t)\right\|^{2}\right].$ (3)
where we align the generated features with the person image, thus encouraging $F_{E}$ attention keys and values to retain valuable information of the garment worn by the model.
Visual-only Modulation Space and Attention. The modulation space of the feature extractor $F_{E}$ follows AdaLN huang2017arbitrary and receives features encoded with CLIP ViT-L radford2021learning and Open-CLIP bigG/14 cherti2023reproducible . For simplicity, we will refer to this operation as $\operatorname{CLIP}(c)$, with $c$ being a textual or visual input. The feature extractor $F_{E}$ does not necessitate text as input, because we want to capture the garment details from $\bm{x}_{\text{model}}$, without “steering” the output process with text. For this reason, we encode the visual projection $\bm{e}^{v}_{\text{pool}}=\operatorname{CLIP}(\bm{x}_{\text{model}})\in\mathbb{R}^{2048}$, which is subsequently used to modulate the latent $\bm{z}_{t}$ from AdaLN estimated scale $\gamma$ and shift $\beta$ as follows:
$\bm{y}_{t}$ $=\operatorname{MLP}(t,\bm{e}^{v}_{\text{pool}})$
$\bm{z}_{t}$ $\leftarrow\gamma(\bm{y}_{t})\bm{z}_{t}+\beta(\bm{y}_{t})$ (4)
using an MLP to jointly encode the pooled vector $\bm{e}^{v}_{\text{pool}}$ together with the timestep $t\in(0,1)$.
From Eq. 4 we notice how AdaLN modulation can shift the distribution $\bm{z}_{t}$ according to “appearance” or “style” of our conditioning pooled input. This justifies our architectural choice, as the latent $\bm{z}_{t}$ in $F_{E}$ inherits the garment information from $\bm{e}^{v}_{\text{pooled}}$ through AdaLN and then propagates it through its layers. Following the reasoning above, we stick to visual-only self-attention modules for $F_{E}$ DiT blocks with queries, keys and values as $\bm{Q}=\bm{Q}_{\bm{z}_{t}},\bm{K}=\bm{K}_{\bm{z}_{t}},\bm{V}=\bm{V}_{\bm{z}_{t}}$. Although this design strategy proves effective for $F_{E}$, we show in Sec. 3.2 that our main DiT $F_{D}$ needs to incorporate three diverse types of input, leading to a new Multimodal Hybrid Attention (MHA) module.
3.2 Text-Enhanced Garment Try-Off
Previous VTON and VTOFF works leverage visual-only inputs, as they claim to have a better transfer of information from the conditioning input to the DiT generation stream. While this approach turns out to be effective for VTON tasks, it has two key shortcomings in the VTOFF setting. First, it prevents our method from generating occluded garments (e.g., a bodysuit worn under a pair of pants). Second, it limits the application of VTOFF to single-category data collections (e.g., upper-body only). Instead, we propose a new, text-enhanced DiT $F_{D}$ to solve these two issues.
Multimodal Hybrid Attention. For the first problem, we assume that it can be addressed by captions focused on structural details and hidden parts. To validate our claim, we propose a new MHA module to seamlessly mix text information, intermediate features from $F_{E}$ and latent features of the denoising DiT. Inspired by the key findings in SD3 esser2024scaling , we concatenate the text features with the visual inputs along the sequence length dimension, thus obtaining:
$\bm{Q}=[\bm{Q}_{\bm{z}_{t}},\bm{Q}_{\text{text}}]\quad\bm{K}=[\bm{K}_{\bm{z}_{t}},\bm{K}_{\text{extractor}},\bm{K}_{\text{text}}]\quad\bm{V}=[\bm{V}_{\bm{z}_{t}},\bm{V}_{\text{extractor}},\bm{V}_{\text{text}}].$ (5)
This module allows the features $\bm{Q}_{\text{text}}$ to attend both the latent projection $\bm{K}_{\bm{z}_{t}}$ and the extractor features $\bm{K}_{\text{extractor}}$. The resulting attention matrix $\bm{A}_{\text{MHA}}$ captures three key interactions: (i) $\bm{A}_{\text{text}\leftrightarrow\bm{z}_{t}}$, preserving pre-trained alignment between language and latent image tokens, (ii) $\bm{A}_{\bm{z}_{t}\leftrightarrow\text{extractor}}$, facilitating transfer between the input garment and the person representation, and (iii) $\bm{A}_{\text{text}\leftrightarrow\text{extractor}}$, grounding the text in the structural features provided by the extractor.
Text embeddings are constructed via the concatenation of CLIP radford2021learning ***As previously mentioned, we consider the combined embedding from CLIP ViT-L and Open-CLIP bigG/14. and T5 raffel2020exploring encoders applied to the input caption $c$ as follows:
$\bm{e}_{\text{text}}=[\operatorname{CLIP}(c),\operatorname{T5}(c)],\quad\text{with }\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}.$ (6)
Visual Category Conditioning with Text Modulation. While some datasets are restricted to upper-body garments choi2021vitonhd , others, such as Dress Code morelli2022dresscode , encompass a broader range, including lower-body items and full-body dresses. This variability introduces ambiguity in garment structure and scale, motivating the need for an explicit encoding of category-level priors.
To address this, we employ high-level conditioning via AdaLN modulation huang2017arbitrary . As shown in previous works garibi2025tokenverse , these layers can be successfully leveraged to adapt “appearance” or “style” information into existing Transformer-based architectures. For this reason, we extract a pooled textual representation $\bm{e}_{\text{pooled}}\in\mathbb{R}^{2048}$ of CLIP textual features of the caption $c$ and inject them into the model through the modulation layers, following Eq. 4. The pooled vector $\bm{e}_{\text{pooled}}\in\mathbb{R}^{2048}$ encapsulates a coarser representation than the full textual embeddings $\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}$, thus being suitable for high-level information conditioning.
We generate textual descriptions using Qwen2.5-VL bai2025qwen2 in a zero-shot setting. We select this model due to its state-of-the-art performance among open-source multimodal large language models. To align with our objective of capturing fine-grained structural attributes, we steer the captioning process to emphasize garment-level semantics, such as garment type, cut, or sleeve length, while deliberately omitting low-level visual features (e.g., color or texture). Empirically, we find that our proposed MHA module is sufficient for accurate color transfer without explicit textual supervision, allowing us to focus the text modality on structural conditioning.
We train the main DiT module (i.e., $F_{D}$) following a diffusion loss with multiple conditioning signals:
$\mathcal{L}_{\text{DiT}}=\mathbb{E}_{\bm{z}_{g},\epsilon_{t},t}\left[\left\|\bm{z}_{g}-F_{D}(\bm{z}_{t},\bm{e}_{\text{pooled}},F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0),t)\right\|^{2}\right],$ (7)
with $F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0)$ being the list of keys and values extracted from $F_{E}$ at timestep $t=0$. We extract this list from $F_{E}$ at $t=0$ and re-use them in $F_{D}$ for all subsequent timesteps, as we want to use key/values from clean data.
3.3 Garment Aligner
While our model is effective at generating realistic and structurally coherent garments, we observe occasional failures in preserving high-frequency details such as fine-grained textures and logos. We hypothesize two primary contributing factors: (i) the diffusion loss $\mathcal{L}_{\text{diff}}$, defined in the noise space, optimizes over perturbed latents rather than directly over image-space reconstructions, limiting its sensitivity to fine-grained patterns; and (ii) the inherent generation dynamics of diffusion models, where errors introduced in early timesteps – typically encoding low-frequency content – can accumulate and degrade the fidelity of high-frequency details in later stages. To mitigate this, we draw inspiration from REPA yu2025representationalignmentgenerationtraining , and propose to explicitly align the internal feature representation of our DiT with that of a pre-trained vision encoder. Specifically, we encourage patch-wise consistency between the eighth Transformer block of our main DiT model $F_{D}$ and the corresponding features extracted from DINOv2 oquab2024dinov2learningrobustvisual .
Let $\bm{h}_{\text{DiT}}\in\mathbb{R}^{3072\times d}$ denote the token sequence obtained from the eighth Transformer block of the DiT decoder $F_{D}$, corresponding to a $64\times 48$ patch grid with embedding dimension $d$. Separately, let $\bm{h}_{\text{enc}}\in\mathbb{R}^{1024\times d^{\prime}}$ be the $32\times 32$ token grid extracted from a frozen DINOv2 encoder, with embedding dimension $d^{\prime}$ (where $d^{\prime}\neq d$). To bridge this mismatch, we introduce a lightweight garment aligner module composed of a convolutional neural network $\phi_{\text{CNN}}:\mathbb{R}^{64\times 48\times d}\rightarrow\mathbb{R}^{32\times 32\times d^{\prime}}$ which is used to downsample the spatial token grid while preserving local structure and to project the token embeddings into the DINOv2 feature space. The aligned tokens are defined as $\tilde{\bm{h}}_{\text{DiT}}=\phi_{\text{CNN}}(\bm{h}_{\text{DiT}})\in\mathbb{R}^{1024\times d^{\prime}}$.
We then enforce feature-level consistency via a cosine similarity loss:
$\mathcal{L}_{\text{align}}=-\mathbb{E}_{\bm{z}_{g},\epsilon_{t},t}\left[\frac{1}{N}\sum_{i=1}^{N}\cos\left(\tilde{\bm{h}}_{i}^{\text{DiT}},\bm{h}_{i}^{\text{enc}}\right)\right],$ (8)
where $\tilde{\bm{h}}_{i}^{\text{DiT}}$ and $\bm{h}_{i}^{\text{enc}}$ are the $i$-th aligned and reference tokens, respectively, $i$ is the patch index, $N$ is the total number of tokens, and $\cos$ is the cosine similarity.
Overall Loss Function. Our final training objective combines the standard diffusion loss $\mathcal{L}_{\text{diff}}$ with the garment alignment loss $\mathcal{L}_{\text{align}}$ previously introduced. The overall objective is thus defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\lambda\cdot\mathcal{L}_{\text{align}},$ (9)
where $\lambda$ is a hyperparameter that balances the contribution of the two loss components.
4 Experiments
4.1 Experimental Settings and Datasets
We conduct our experiments using two publicly available fashion datasets: VITON-HD choi2021vitonhd and Dress Code morelli2022dresscode . Both datasets provide images at a resolution of $1024\times 768$. VITON-HD contains only upper-body garments and represents a single-category setting, while Dress Code includes multiple categories (i.e., dresses, upper-body, and lower-body garments) enabling evaluation of the generalization capabilities of our methods across diverse garment types. For both the feature extractor and the diffusion backbone, we adopt Stable Diffusion 3 medium esser2024scaling . All models are trained on a single node equipped with 4 NVIDIA A100 GPUs (64GB each), using DeepSpeed ZeRO-2 rajbhandari2020zeromemoryoptimizationstraining for efficient distributed training. We use a total batch size of 32 and train each model for 30k steps, corresponding to approximately 960k images. Optimization is performed with AdamW loshchilov2019decoupledweightdecayregularization , using a learning rate of $1\times 10^{-4}$, a warmup phase of 3k steps, and a cosine annealing schedule. We train separate models per dataset to account for differences in distribution and garment structure. In all experiments, we set the alignment loss weight $\lambda$ equal to $0.5$.
Table 1: Quantitative results on the Dress Code dataset, considering both the entire test set and the three category-specific subsets. $\uparrow$ indicates higher is better, $\downarrow$ lower is better.
All Upper-Body
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 75.79 39.70 29.88 70.02 42.80 76.59 40.62 29.04 37.97 17.30
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 77.56 35.17 25.17 12.32 3.65 76.61 38.99 25.78 15.77 3.22
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 74.54 35.48 19.75 10.94 0.76
Lower-Body Dresses
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 74.10 43.35 32.54 162.02 137.10 76.79 35.29 27.88 72.19 47.58
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 78.15 34.72 25.87 30.06 12.01 77.93 31.80 23.86 19.20 6.27
TEMU-VTOFF (Ours) 73.94 34.60 19.57 13.83 2.04 79.39 24.32 16.67 11.29 0.59
Figure 3: Qualitative comparison on the Dress Code dataset between images generated by TEMU-VTOFF and those generated by competitors.
4.2 Comparison with the State of the Art
To evaluate the proposed TEMU-VTOFF architecture, we use a combination of perceptual, structural, and distributional similarity metrics. Specifically, we report LPIPS zhang2018perceptual , SSIM wang2004ssim , DISTS Ding_2020 , FID parmar2021cleanfid , and KID bińkowski2021demystifyingmmdgans . We compare our approach against recent VTOFF methods, including TryOffDiff velioglu2024tryoffdiff , TryOffAnyone xarchakos2024tryoffanyone , and Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition , the latter being a more general framework designed for both VTOFF and VTON tasks. Notably, TryOffDiff and TryOffAnyone are trained exclusively on VITON-HD, which contains only upper-body garments. In contrast, Any2AnyTryon is trained on a mixture of datasets including Dress Code, VITON-HD, and DeepFashion2 ge2019deepfashion2versatilebenchmarkdetection , enabling broader category coverage.
For fair comparison on the multi-category Dress Code dataset, we retrain TryOffDiff from scratch using its publicly available code and default settings. Due to the lack of released training code, we are unable to retrain TryOffAnyone. This setup allows us to properly evaluate how well existing methods generalize across multiple categories and to highlight the robustness and flexibility of our approach in handling different garment types.
Results on the Dress Code Dataset. Table 1 presents the image quality metrics on the Dress Code dataset. Our method significantly outperforms existing state-of-the-art approaches across all evaluation metrics. The results are especially pronounced in this dataset, with respect to prior works, because of its multi-category setting. This shows that our method is category-agnostic and benefits from the joint use of textual garment descriptions and fine-grained masks. As a result, our model achieves a better perceptual quality and better alignment with the ground-truth distribution.
In Fig. 3, we provide qualitative results comparing TEMU-VTOFF with competitors. This shows the challenges introduced by the Dress Code different set of categories. TryOffDiff often fails to preserve key visual characteristics such as color, texture and shape. Furthermore, it struggles with the reconstruction of lower-body items. Instead, Any2AnyTryon performs better, but often fails with the generation of the target garment. In contrast, our method is able to closely match the target garment across all categories.
Results on the VITON-HD Dataset. In Table 2, we report the quantitative results on the VITON-HD dataset. Also in this setting, TEMU-VTOFF achieves consistent improvements over existing approaches across all metrics. However, the improvements are less pronounced compared to those observed on the Dress Code dataset. This discrepancy is primarily due to the more constrained nature of VITON-HD, which focuses exclusively on upper-body garments. By contrast, Dress Code includes different categories, including full-body and lower-body items like dresses, skirts, and pants, which are more prone to occlusions and complex geometries. In these cases, the use of textual descriptions and fine-grained masks becomes crucial, offering strong guidance that significantly enhances garment reconstruction quality. Thus, the improvements of our approach are especially evident in more challenging multi-category scenarios.
A visual comparison on sample VITON-HD images is shown in Fig. 4, which further demonstrates the improved garment reconstruction quality of our proposed method.
Table 2: Quantitative results on the VITON-HD dataset. $\uparrow$ indicates higher is better, $\downarrow$ lower is better.
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 75.53 39.56 25.53 17.49 5.30
TryOffAnyone xarchakos2024tryoffanyone 75.90 35.26 23.47 12.74 2.85
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 75.72 37.95 24.32 12.88 3.01
TEMU-VTOFF (Ours) 77.21 28.44 18.04 8.71 1.11
Figure 4: Qualitative comparison on the VITON-HD dataset between images generated by TEMU-VTOFF and those generated by competitors.
4.3 Ablation Studies
To assess the contribution of each component in our pipeline, we conduct a detailed ablation study on the Dress Code dataset, as shown in Table 3. Specifically, we evaluate the impact of removing the textual description of the garment, the fine-grained mask, or both. Further, we investigate the impact of our garment aligner. The results show that all the components play a crucial role, demonstrating their contribution to improving the final image quality.
Table 3: Ablation study of the proposed components on the Dress Code dataset.
All Upper-body Lower-body Dresses
SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$
Effect of Text and Mask Conditioning
w/o text and masks 71.04 39.68 25.20 9.63 3.17 23.71 19.75 65.85 49.19 20.12 15.47
w/o text modulation 73.88 34.63 22.54 7.75 1.52 24.02 13.48 24.33 18.13 19.27 13.30
w/o fine-grained masks 74.65 32.33 20.87 6.58 1.03 20.85 11.31 22.34 15.74 19.42 13.62
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
Effect of Garment Aligner Component
w/o garment aligner 76.01 30.84 20.63 5.91 0.78 21.77 11.26 22.26 14.22 17.86 11.86
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
(a) Evaluation of mask and text joint impact.
(b) Evaluation of garment aligner impact.
Figure 5: Qualitative comparisons validating the effectiveness of the proposed components on the Dress Code dataset.
To better understand the strength of each component proposed with our approach, in Fig. 5 we provide a visual comparison on the Dress Code dataset. Our method without textual conditioning relies exclusively on visual features extracted from the person wearing the garment, without any textual guidance. Notably, this often results in failures in challenging cases, especially for garments with strong occlusions such as bodysuits or garments partially hidden by arms or hair. This limitation has been addressed by conditioning the generation to also use a textual structural description of the target garment. This structural guidance enables the model to better capture the structure of the garment, leading to significantly more accurate outputs in occluded scenarios. In addition, using fine-grained masks and the garment aligner further improves the quality of the generated garments.
Limitations. While our method shows strong performance and generalization, some limitations remain. First, it struggles to reconstruct fine-grained details such as logos or printed text – partly due to the SD3 medium backbone, which is known to handle such elements inconsistently. Second, performance on lower-body garments is less reliable than for upper-body garments and dresses, likely due to class imbalance in the Dress Code dataset. Additional discussion and failure cases are included in the supplementary material.
5 Conclusion
We have presented TEMU-VTOFF, a novel virtual try-off architecture designed to generate high-quality, diverse in-shop garment images. Our approach is the first to effectively extend VTOFF to multi-category scenarios, addressing existing inherent shortcomings of existing methods with ad-hoc modules such as a dual-DiT architecture equipped with our multimodal hybrid attention to effortlessly integrate garment, person, and text features. Moreover, we have proposed an external garment aligner module paired with a new loss objective for visual detail refinement. Experimental results on standard VTOFF benchmarks demonstrate the robustness and effectiveness of our approach across a range of garment types and evaluation metrics.
[…]
Appendix A Experimental Protocols
A.1 Datasets Details
Dress Code. In our experiments, we adopt the Dress Code dataset [12], the largest publicly available benchmark for image-based virtual try-on. Unlike previous datasets limited to upper-body clothing, Dress Code includes three macro-categories:
  • Upper-body: $15,363$ pairs (e.g., tops, t-shirts, shirts, sweatshirts)
  • Lower-body: $8,951$ pairs (e.g., trousers, skirts, shorts)
  • Dresses: $29,478$ (e.g., full body dresses)
The total number of paired samples is $53,792$, split into $48,392$ training images and $5,400$ test images at a resolution of $1024\times 768$.
VITON-HD. Following previous literature, we also adopt VITON-HD [8], a publicly available dataset widely used in virtual try-on research. It is composed exclusively of upper-body garments and provides high-resolution images at $1024\times 768$ pixels. The dataset contains a total of $27,358$ images, structured into $13,679$ garment-model pairs. These are split into $11,647$ training pairs and $2,032$ test pairs, each comprising a front-view image of a garment and the corresponding image of a model wearing it.
A.2 Implementation Details
We evaluate our method both with distribution-based metrics and per-sample similarity metrics. For the first group, we adopt FID [37] and KID [4] implementations derived from clean-fid PyTorch packagehttps://pypi.org/project/clean-fid/. Concerning the second group, we adopt both SSIM [50] and LPIPS [58] as they are the standard metrics adopted in the field to measure structural and perceptual similarity between a pair of images. We reuse the corresponding Python packages provided by TorchMetricshttps://pypi.org/project/torchmetrics/. Finally, we adopt DISTS [13] as an additional sample-based similarity metric, as it correlates better with human judgment, as shown in previous works [16]. We stick to the corresponding Python package§§§https://pypi.org/project/DISTS-pytorch/ to compute it for our experiments.
A.3 Algorithm
To provide a clear understanding of TEMU-VTOFF, we summarize the core components of our method in Algorithm 1. The pseudo-code outlines the sequential steps involved in training our dual-DiT architecture, including multimodal conditioning, the hybrid attention module, and the garment aligner component.
Algorithm 1 TEMU-VTOFF: Virtual Try-Off with Dual-DiT and Garment Alignment
1:Person image $\bm{x}_{\text{model}}$, garment caption $c$, binary mask $M$, target garment image $\bm{x}_{g}$
2:Generated garment $\hat{\bm{x}}_{g}$
3:Latent encoding:
4: Encode the target garment: $\bm{z}_{g}\leftarrow\mathcal{E}(\bm{x}_{g})$
5: Sample noise: $\epsilon_{t}\sim\mathcal{N}(0,1)$
6: Apply flow-matching: $\bm{z}_{t}\leftarrow(1-t)\bm{z}_{g}+t\cdot\epsilon_{t}$
7:Prepare masked spatial input:
8: Encode masked person image: $\bm{x}_{M}\leftarrow\mathcal{E}(\bm{x}_{\text{model}}\odot M)$
9: Concatenate inputs: $\bm{z}^{\prime}_{t}\leftarrow[\bm{z}_{t},M,\bm{x}_{M}]$
10:Extract modulation features:
11:$\bm{e}^{v}_{\text{pool}}\leftarrow\text{CLIP}(\bm{x}_{\text{model}})$
12:Extract keys and values using feature extractor:
13:$\{\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}\}_{l=1}^{N}\leftarrow F_{E}(\bm{z}^{\prime}_{0},\bm{x}_{\text{model}},t{=}0)$
14:Encode text information:
15: Get pooled text embedding: $\bm{e}_{\text{pooled}}\leftarrow\text{CLIP}(c)$
16: Get full sequence text features: $\bm{e}_{\text{text}}\leftarrow[\text{CLIP}(c),\text{T5}(c)]$
17:Noise prediction:
18:$\hat{\epsilon}_{t}\leftarrow F_{D}(\bm{z}_{t},\bm{e}_{\text{pooled}},\bm{e}_{\text{text}},\{\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}\},t)$
19: Compute diffusion loss: $\mathcal{L}_{\text{diff}}\leftarrow\|\hat{\epsilon}_{t}-\epsilon_{t}\|^{2}$
20:Align internal representations:
21: Extract DiT features: $\bm{h}_{\text{DiT}}\leftarrow$ tokens from 8th block of $F_{D}$
22: Extract DINOv2 features: $\bm{h}_{\text{enc}}\leftarrow\text{DINOv2}(\bm{x}_{g})$
23: Align via projection: $\tilde{\bm{h}}_{\text{DiT}}\leftarrow\phi_{\text{CNN}}(\bm{h}_{\text{DiT}})$
24: Compute alignment loss: $\mathcal{L}_{\text{align}}\leftarrow-\frac{1}{N}\sum_{i}\cos(\tilde{\bm{h}}_{i}^{\text{DiT}},\bm{h}_{i}^{\text{enc}})$
25:Final objective:
26: Combine losses: $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{diff}}+\lambda\cdot\mathcal{L}_{\text{align}}$
27:Decode final garment:
28: Run reverse process: $\hat{\bm{x}}_{g}\leftarrow\mathcal{D}(\hat{\bm{z}}_{0})$
Appendix B Caption Extraction Details
We leverage Qwen2.5-VL [1] to generate the caption of a given garment image, following the chat template provided below:
visual_attributes = {
"dresses": ["Cloth Type", "Waist", "Fit", "Hem", "Neckline", "Sleeve Length", "Cloth Length"],
"upper_body": ["Cloth Type", "Waist", "Fit", "Hem", "Neckline", "Sleeve Length", "Cloth Length"],
"lower_body": ["Cloth Type", "Waist", "Fit", "Cloth Length"]
category = "upper_body" # The category to which the image belongs
System: You are Qwen, created by Alibaba Cloud. You are a helpful assistant.
User:[Uncaptioned image] Use only visual attributes that are present in the image. Predict values of the following attributes: {visual_attributes[category]}. It’s forbidden to generate the following visual attributes: colors, background, and textures/patterns. It’s forbidden to generate unspecified predictions. It’s forbidden to generate newline characters. Generate in this way: a <cloth type> with <attributes description>
Qwen Caption: A denim shirt with a straight fit, long sleeves, and a button-down neckline. The hem is straight and the shirt appears to be of standard length.
We decide to generate structural-only attributes because our base model without text can already transfer colors and textures correctly from the person image to the generated garment image. The structural attributes are slightly different according to the three categories of clothing, as specified in visual_attributes. For example, the neckline can be specified for upper body and dresses (whole body garments), but not for lower body items.
Appendix C Additional Qualitative Results
We show an extended version of the qualitative results presented in our main paper. We report additional visual results on the competitors in Fig. F on the Dress Code [12] dataset and in Fig. G on the VITON-HD [8] dataset. Furthermore, Fig. H presents additional ablation results to analyze the impact of textual and mask conditioning. Finally, we include in Fig. I the full set of inputs used for generating the target garment, including the model input, the segmentation mask, and the caption.
Figure F: Additional qualitative results of TEMU-VTOFF and competitors on Dress Code [12].
Figure G: Additional qualitative results of TEMU-VTOFF and competitors on VITON-HD [8].
Figure H: Additional qualitative results showing the contribution of each component in TEMU-VTOFF on Dress Code [12] images.
Figure I: Inputs used to generate the target garment with TEMU-VTOFF, using sample images from Dress Code [11]
Appendix D Limitations
Our model inherits some inner problems of foundational models like Stable Diffusion 3 [14]. Even though our method improved the generation of big logos and text, it is limited in its scope concerning fine-grained details like complex texture patterns and small written text. Moreover, it sometimes fails to render the correct number of small objects like buttons. We show a set of failure cases in Fig. J and Fig. K, on sample images from Dress Code [11] and VITON-HD [32] respectively.
Appendix E Broader Impact
Our method addresses the VTOFF task by generating flat, in-shop garment images from photos of dressed individuals. This enables a novel form of data augmentation in the fashion domain, allowing clean garment representations to be synthesized without manual segmentation or dedicated photoshoots. By bridging the gap between worn and catalog-like appearances, our approach can improve scalability for fashion datasets and support downstream applications such as retrieval, recommendation, and virtual try-on.
However, as with any generative technology, there are important ethical and legal considerations. In particular, our model could be used to reconstruct garments originally designed by third parties, potentially raising issues of copyright and intellectual property infringement. We emphasize that our framework is intended for research and responsible use, and any deployment in commercial settings should ensure compliance with applicable copyright laws and respect for designer rights.
Figure J: An overview of failure cases on the Dress Code [12] dataset.
Figure K: An overview of failure cases on the VITON-HD [8] dataset.